# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ 6 Conclusion
We revisit imitation learning from the perspective of state-wise annotation. We show via the Stagger algorithm that, interaction with the demonstrating expert, with its cost properly measured, can enable provable cost efficiency gains over Behavior Cloning. We also propose Warm-Stagger that combines the benefits of offline data and interactive feedback. Our theoretical construction shows that such a hybrid method can strictly outperform both pure offline and pure interactive baselines under realistic cost models. Empirical results on the synthetic MDP support our theoretical findings, while MuJoCo experiments demonstrate the practical viability and competitive performance of our methods on continuous control tasks. Additionally, we show a trajectory-wise annotation variant of DAgger can match the sample complexity of log-loss Behavior Cloning without recoverability assumptions (Appendix E), with additional experiments (Appendix G.3).
@@
Our design of imitation learning algorithm only aims at closing the gap between the performance of the expert and the trained policy; thus, the performance of our learned policy is bottlenecked by the expert’s performance. In this respect, designing imitation learners outputting policies that surpass expert performance is an important direction.
Our theory provide sample complexity guarantees for the discrete-action setting with deterministic and realizable expert. When such assumptions are relaxed, additional challenges arise simchowitz2025pitfalls. In this respect, there remains a gap between our theoretical analysis and our MuJoCo experiment results. In future work, we are interested in conducting additional experiments on discrete-action control problems (e.g., Atari) as well as language model distillation tasks.
Acknowledgments.
We thank the anonymous NeurIPS reviewers for their helpful feedback, which significantly improved the presentation of the paper. We thank National Science Foundation IIS-2440266 (CAREER) for research support.
Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning — Supplementary Materials
@@ Appendix A Additional Related Work
The emergence of first-step mixing policies originated from technical considerations. In may interactive IL methods [ross2011reduction, ross2014reinforcement], the returned policy was not a uniform first-step mixture but rather the best policy selected through validation. However, performing such validation in an interactive setting often requires additional expert annotations. Subsequent works [rajaraman2021value, li2022efficient, li2023agnostic, foster2024behavior] circumvented the need for validation by employing a uniform first-step mixture of policies across learning rounds, thereby directly translating online regret guarantees into performance differences. Our Tragger algorithm (Algorithm 3 in Appendix E) also employs a first-step mixing policy at each iteration, and has sample complexity on par with behavior cloning.
On the other hand, each-step mixing between the learned policy across rounds and the expert policy has been a prevalent strategy in interactive IL approaches [daume2009search, ross2010efficient, ross2011reduction, ross2014reinforcement]. For each-step mixture policies, [li2022efficient] was the first to explicitly distinguish this approach from first-step mixing. In other works [rajaraman2021value, foster2024behavior], each-step mixing can be interpreted as learning $H$ separate mixture policies, one for each step within an episode.
@@
Though this work follows [foster2024behavior] and focuses on log loss, we believe the $1/n$-rate is not exclusive to log loss. Despite requiring an additional supervision oracle, [le2018hierarchical] suggests that trajectory-wise annotation complexity similar to Theorem 3 (and Theorem 27) can be achieved using Halving [shalev2011online] and 0-11 loss.
From an algorithmic perspective, we explored trajectory-wise annotation with first-step mixing (Algorithm 3 in Appendix E) and state-wise annotation with each-step mixing (Algorithm 1). For trajectory-wise annotation with each-step mixing, naively learning a parameter-sharing policy may encounter a batch-summed log loss, introducing an (undesirable) additional $H$ factor to the sample complexity. Analyzing state-wise annotation with each-step mixing remains an open question for future research.
In terms of practical implementation, it is worth noting that even with oracle-efficient implementations (e.g. [li2022efficient, li2023agnostic]), interactive IL may require multiple supervised learning oracle calls per iteration. In contrast, offline IL requires only a single oracle call to obtain the returned policy, which provides a clear computational advantage. We also note that real-world experts can be suboptimal; in some applications it may be preferable to combine imitation and reinforcement learning signals (e.g., [ross2014reinforcement, sun2018truncated, amortila2022few]).
Lower bounds in interactive imitation learning.
From an information-theoretic perspective, a line of work [rajaraman2020toward, rajaraman2021value, foster2024behavior] provides lower bounds for imitation learning under the realizable setting and considers $\mu$-recoverability. [rajaraman2021value] is the first to demonstrate a gap between the lower bounds of offline IL and interactive IL in trajectory-wise annotation, focusing on the tabular and non-parameter-sharing setting. [foster2024behavior] establishes a $\Omega\left(\frac{H}{\epsilon}\right)$ sample-complexity lower bound for trajectory-wise annotation in the parameter-sharing setting.
We observe that the proof of  [foster2024behavior, Theorem 2.2] also implicitly implies a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for the state-wise annotation setting. Their proof relies on an MDP consisting only of self-absorbing states, where annotating a full trajectory gives the same amount of information as annotating a single state. In that MDP (which is $1$-recoverable), Algorithm 1 achieves $\tilde{O}(\frac{H\log(B)}{\epsilon})$ state-wise sample complexity, which does not contradict this lower bound. Nonetheless, obtaining lower bounds for state-wise sample complexity for general MDPs, policy classes, and general recoverability constants remains an open question.
@@ Appendix B Proof for Stagger
We first present 2 useful distance measures for pair of policies.
@@
$\lambda(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(\cdot|s_{1:H})}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq a^{\prime}_{h})\right].$
$\lambda(\pi\parallel\pi^{\prime})$ is the expected total number of actions taken by $\pi^{\prime}$ that deviates from actions in trajectories induced by $\pi$. Note that $\lambda(\cdot||\cdot)$ is asymmetric, while the same concept is applied in offline and interactive IL [ross2010efficient, ross2011reduction] with different guarantees for $\lambda(\pi\parallel\pi^{\mathrm{E}})$ and $\lambda(\pi^{\mathrm{E}}\parallel\pi)$ (Lemma 36).
@@
State-wise Hellinger distance represents the expected Hellinger distance between the action distribution of $\pi$ and $\pi^{\prime}$ on $s\sim d^{\pi}$. One notable feature here is that the distance is evaluated between $\pi(\cdot\mid s)$ and $\pi^{\prime}(\cdot\mid s)$, independent of the original action $a$ taken by $\pi$ when visiting $s$. By Lemma 36, state-wise Hellinger distance can relate to trajectory-wise $L_{1}$-divergence.
In the following, we show that the performance difference between the policy $\hat{\pi}$ returned by Stagger (Algorithm 1) and the expert policy $\pi^{\mathrm{E}}$ can be bounded by the state-wise Hellinger estimation error:
@@
where $\pi^{n}(\cdot\mid s)$ and $\pi^{E}_{h}(\cdot\mid s)$ denote the action distributions over $\mathcal{A}$ produced by the policies $\pi^{n}$ and $\pi^{E}$ at state $s$.
We first prove this in Lemma 10, and then prove the state-wise annotation complexity of Algorithm 1 in Theorem 11.
Different from [foster2024behavior], where access to full action demonstrations is assumed, we consider a more restrictive model where, at each round $n$, only a single state $(s^{n},h^{n})$ from the trajectory induced by $\pi^{n}$ is sampled and annotated by the expert.
@@
$\lambda(\pi\parallel\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq\pi^{\mathrm{E}}(s_{h}))\right]=\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
@@
Observe that $H\cdot\mathbb{E}_{s\sim d^{\pi}}\left[D_{H}^{2}(\pi(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right]=\sum_{h=1}^{H}\mathbb{E}^{\pi}\left[D_{H}^{2}(\pi(\cdot\mid s_{h}),\pi^{\mathrm{E}}(\cdot\mid s_{h}))\right]$, we conclude the proof by applying Lemma 35 with $p=\pi(s_{h})$ and $q=\pi^{\mathrm{E}}(s_{h})$, which gives
@@
Recall the each-step mixing in Definition 1, since $\bar{\pi}_{u}$ is a each-step mixing policy, where $\forall h\in[H],s\in{\mathcal{S}}$, $\bar{\pi}_{u}(a|s)=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s)$.
By using $\bar{\pi}_{u}$, the loss functions at each round that passed through online learning oracle $\mathbb{A}$, $\ell^{n}(\pi)$ is of the form
$\ell^{n}(\pi)=\log(1/\bar{\pi}_{u}(a^{n,*}|s^{n}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}|s^{n})}\right),$
@@
$\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{n}_{h^{n}}(a^{*,n}\mid s^{n}))\leq\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{\mathrm{E}}_{h^{n}}(a^{*,n}\mid s^{n}))+\log(B)=\log(B).$
Then, Lemma 39, a standard online-to-batch conversion argument with $x^{n}=(s^{n},h^{n})$, $y^{n}=a^{*,n}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s^{n},h^{n},a^{n},a^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N_{\mathrm{int}}}=\sum_{n=1}^{{N_{\mathrm{int}}}}\mathbb{E}_{(s^{n},h^{n})\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h^{n}}(\cdot\mid s^{n}),\pi^{\mathrm{E}}_{h^{n}}(\cdot\mid s^{n}))\right]\leq\log(B)+2\log(1/\delta).$
@@ Appendix C Proof for Warm-Stagger
In this section, we analyze the guarantees of Warm-Stagger under the realizable and deterministic expert assumption. We show that all intermediate policies, as well as the final returned mixture policy, enjoy small Hellinger distance to the expert’s trajectory distribution, due to their agreement on the offline dataset. Our analysis builds on maximum likelihood estimator (MLE) generalization guarantees under log-loss minimization, and leverages the fact that each each-step mixing policy in Warm-Stagger can be viewed as a first-step policy mixture. The following claim shows that any Markovian policy $\pi$ can be represented as a first-step mixing over deterministic Markovian policies from a finite class.
@@
Let $\bar{\pi}_{u}$ be a each-step mixing policy that, at each step $h\in[H]$, samples a base policy $\pi\in\mathcal{B}$ independently according to $u\in\Delta(\mathcal{B})$. Then, its induced trajectory distribution is equivalent to a first-step mixing the each-step policy completion of $\mathcal{B}$, denoted by $\tilde{\mathcal{B}}:=\left\{(\pi_{1},\dots,\pi):\pi\in\mathcal{B}\right\}$ (see Definition 4).
@@
Let $\nu=(\pi_{1},\dots,\pi_{H})\in\tilde{\mathcal{B}}$. Define $u(\nu):=\prod_{h=1}^{H}u(\pi_{h})$, which is a valid distribution over $\tilde{\mathcal{B}}$. Consider the joint action distribution under $\bar{\pi}_{u}$, which samples $\pi_{h}\sim u$ independently for each step and executes $a_{h}\sim\pi_{h}(\cdot\mid s_{h})$. The resulting conditional distribution over actions given the state sequence is
$\bar{\pi}_{u}(a_{1:H}\mid s_{1:H})=\prod_{h=1}^{H}\left(\sum_{\pi_{h}\in\mathcal{B}}u(\pi_{h})\,\pi_{h}(a_{h}\mid s_{h})\right).$
Under the first-step mixture policy $\pi_{u(\nu)}$ over $\tilde{\mathcal{B}}$, a full tuple $\nu=(\pi_{1},\dots,\pi_{H})$ is sampled once from $u(\nu)$, and actions are drawn as $a_{h}\sim\pi_{h}(\cdot\mid s_{h})$. The resulting action distribution is
$\pi_{u(\nu)}(a_{1:H}\mid s_{1:H})=\sum_{\nu\in\tilde{\mathcal{B}}}u(\nu)\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
@@
$\sum_{(\pi_{1},\dots,\pi_{H})}\left(\prod_{h=1}^{H}u(\pi_{h})\pi_{h}(a_{h}\mid s_{h})\right)=\prod_{h=1}^{H}\sum_{\pi_{h}\in\mathcal{B}}u(\pi_{h})\,\pi_{h}(a_{h}\mid s_{h}),$

### v3

@@ 6 Conclusion
We revisit imitation learning from the perspective of state-wise annotations. We show via the Stagger algorithm that, interaction with the demonstrating expert, with its cost properly measured, can enable provable cost efficiency gains over Behavior Cloning. We also propose Warm-Stagger that combines the benefits of offline data and interactive feedback. Our theoretical analysis shows that such a hybrid method can strictly outperform both pure offline and pure interactive baselines under realistic cost models. Empirical results on our synthetic MDP support our theoretical findings, while MuJoCo experiments demonstrate the practical viability and competitive performance of our methods on continuous control tasks. In Appendix E, we also show that a trajectory-wise annotation variant of DAgger can match the sample complexity of log-loss Behavior Cloning without recoverability assumptions, with additional experiments (Appendix G.3).
@@
Our design of imitation learning algorithm only aims at closing the gap between the performance of the expert and the trained policy; thus, the performance of our learned policy is bottlenecked by the expert’s performance. In this respect, designing imitation learners that output policies surpassing expert performance is an important direction.
Our theory provide sample complexity guarantees for the discrete-action setting with deterministic and realizable expert. When such assumptions are relaxed, additional challenges arise Simchowitz et al. (2025). In this respect, there remains a gap between our theoretical analysis and our MuJoCo experiment results. In future work, we are interested in conducting additional experiments on discrete-action control problems (e.g., Atari) as well as language model distillation tasks.
Acknowledgments:
We thank the anonymous NeurIPS reviewers for their helpful feedback, which significantly improved the presentation of the paper. We thank Kianté Brantley for helpful discussions on state-wise annotations. We thank National Science Foundation IIS-2440266 (CAREER) for research support.
@@ Appendix A Additional Related Work
The emergence of first-step mixing policies originated from technical considerations. In may interactive IL methods [51, 50], the returned policy was not a uniform first-step mixture but rather the best policy selected through validation. However, performing such validation in an interactive setting often requires additional expert annotations. Subsequent works [46, 33, 34, 17] circumvented the need for validation by employing a uniform first-step mixture of policies across learning rounds, thereby directly translating online regret guarantees into performance differences. Our Tragger algorithm (Algorithm 3 in Appendix E) also employs a first-step mixing policy at each iteration, and has state-wise sample complexity on par with behavior cloning.
On the other hand, each-step mixing between the learned policy across rounds and the expert policy has been a prevalent strategy in interactive IL approaches [14, 49, 51, 50]. For each-step mixture policies, [33] was the first to explicitly distinguish this approach from first-step mixing. In other works [46, 17], each-step mixing can be interpreted as learning $H$ separate mixture policies, one for each step within an episode.
@@
Though this work follows [17] and focuses on log loss, we believe our $O(1/N)$ rate is not exclusive to log loss. Despite requiring an additional supervision oracle, [31] suggests that trajectory-wise annotation complexity similar to Theorem 3 (and Theorem 31) can be achieved using Halving [57] and 0-11 loss.
From an algorithmic perspective, we explored trajectory-wise annotation with first-step mixing (Algorithm 3 in Appendix E) and state-wise annotation with each-step mixing (Algorithm 1). For trajectory-wise annotation with each-step mixing, naively learning a parameter-sharing policy may encounter a batch-summed log loss, introducing an (undesirable) additional $H$ factor to the sample complexity [21, 74]. Analyzing state-wise annotation with first-step mixing remains an open question.
For practical implementations, it is worth noting that even with oracle-efficient implementations (e.g. [33, 34]), interactive IL may require multiple supervised learning oracle calls per iteration. In contrast, offline IL requires only a single oracle call to obtain the returned policy, which provides a clear computational advantage. We also note that real-world experts can be suboptimal; in some applications it may be preferable to combine imitation and reinforcement learning signals (e.g., [50, 65, 3]).
Information-theoretic lower bounds for interactive imitation learning.
A line of works [47, 46, 17] provides lower bounds for the sample complexity of imitation learning under the realizable setting and considers $\mu$-recoverability. [46] is the first to demonstrate a gap between the lower bounds of offline IL and interactive IL in trajectory-wise annotation, focusing on the tabular and non-parameter-sharing setting. [17] establishes a $\Omega\left(\frac{H}{\epsilon}\right)$ sample-complexity lower bound for trajectory-wise annotation in the parameter-sharing setting.
We observe that the proof of  [17, Theorem 2.2] also implicitly implies a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for the state-wise annotation setting. Their proof relies on an MDP consisting only of self-absorbing states, where annotating a full trajectory gives the same amount of information as annotating a single state. In that MDP (which is $1$-recoverable), Algorithm 1 achieves $\tilde{O}(\frac{H\log(B)}{\epsilon})$ state-wise sample complexity, which does not contradict this lower bound. Nonetheless, obtaining lower bounds for state-wise sample complexity for general MDPs, policy classes, and general recoverability constants remains an open question.
@@ Appendix B Proof for Stagger
We first present two useful distance measures for pair of policies.
@@
$\lambda(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1}\sim\pi^{\prime}(\cdot|s_{1}),\ldots,a^{\prime}_{H}\sim\pi^{\prime}(\cdot|s_{H})}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq a^{\prime}_{h})\right].$
$\lambda(\pi\parallel\pi^{\prime})$ is the expected total number of actions taken by $\pi^{\prime}$ that deviates from actions in trajectories induced by $\pi$. Note that $\lambda(\cdot||\cdot)$ is asymmetric, while the same concept is applied in offline and interactive IL [49, 51] with different guarantees for $\lambda(\pi\parallel\pi^{\mathrm{E}})$ and $\lambda(\pi^{\mathrm{E}}\parallel\pi)$ (see Lemma 39).
@@
State-wise Hellinger distance represents the expected Hellinger distance between the action distribution of $\pi$ and $\pi^{\prime}$ on $s\sim d^{\pi}$. One notable feature here is that the distance is evaluated between $\pi(\cdot\mid s)$ and $\pi^{\prime}(\cdot\mid s)$, unrelated to the original action $a$ taken by by $\pi$ when visiting $s$. By Lemma 38, state-wise Hellinger distance with the expert policy is a constant factor equivalent to trajectory-wise $L_{1}$-divergence in the deterministic realizable expert setting.
In the following lemma, we show that the performance difference between the policy $\hat{\pi}$ returned by Stagger (Algorithm 1) and the expert policy $\pi^{\mathrm{E}}$ can be bounded by the online state-wise Hellinger estimation error:
@@
where $\pi^{n}(\cdot\mid s)$ and $\pi^{E}(\cdot\mid s)$ denote the action distributions produced by the policies $\pi^{n}$ and $\pi^{E}$ at state $s$.
We are ready to prove the state-wise annotation complexity of Algorithm 1:
@@
$\lambda(\pi\parallel\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq\pi^{\mathrm{E}}(s_{h}))\right]=\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi(\cdot\mid s_{h})-\pi^{\mathrm{E}}(\cdot\mid s_{h})\|_{1}.$
@@
Observe that $H\cdot\mathbb{E}_{s\sim d^{\pi}}\left[D_{H}^{2}(\pi(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right]=\sum_{h=1}^{H}\mathbb{E}^{\pi}\left[D_{H}^{2}(\pi(\cdot\mid s_{h}),\pi^{\mathrm{E}}(\cdot\mid s_{h}))\right]$, we conclude the proof by applying Lemma 38 with $p=\pi(\cdot\mid s_{h})$ and $q=\pi^{\mathrm{E}}(\cdot\mid s_{h})$, which gives
@@
Recall the each-step mixing in Definition 1: since $\bar{\pi}_{u}$ is a each-step mixing policy, we have that $\forall h\in[H],s\in{\mathcal{S}}$, $\bar{\pi}_{u}(a|s)=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s)$.
The loss function at each round that passed through online learning oracle $\mathbb{A}$, evaluated at $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$, $\ell^{n}(\bar{\pi}_{u})$, is of the form
$\ell^{n}(\bar{\pi}_{u})=\log\left(\frac{1}{\bar{\pi}_{u}(a^{n,*}|s^{n})}\right)=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}|s^{n})}\right),$
@@
$\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{n}(a^{*,n}\mid s^{n}))\leq\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{\mathrm{E}}(a^{*,n}\mid s^{n}))+\log(B)=\log(B).$
Then, Lemma 42, a standard online-to-batch conversion argument with $x^{n}=(s^{n},h^{n})$, $y^{n}=a^{*,n}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s^{n},a^{n},a^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N_{\mathrm{int}}}=\sum_{n=1}^{{N_{\mathrm{int}}}}\mathbb{E}_{s^{n}\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}),\pi^{\mathrm{E}}(\cdot\mid s^{n}))\right]\leq\log(B)+2\log(1/\delta).$
@@ Appendix C Proof for Warm-Stagger
In this section, we analyze the guarantees of Warm-Stagger under the realizable and deterministic expert assumption. We show that all intermediate policies, as well as the final returned mixture policy, induces distributions over trajectories that enjoy small Hellinger distance to the expert’s trajectory distribution, due to their agreement on the offline dataset. Our analysis builds on generalization guarantees of the maximum likelihood estimator (MLE).
C.1 Preliminaries: First-step mixing and causally conditioned probabilities
Our main analysis leverages the fact that the each-step mixing policies $\pi^{n}$’s maintained by Warm-Stagger can be viewed as first-step mixtures of policies in $\tilde{\mathcal{B}}$ (recall Definition 5) – this subsection is dedicated to prove this result (Lemma 16). We first recall the definition of first-step mixing of policies, adapted to our notations:
Definition 13 (First-step mixing, e.g. [69]).
Given a distribution $u$ over a set of (possibly nonstationary) policies $\Pi$, denote $\pi_{u}$ as the first-step mixing policy induced by $u$: when rolling out $\pi_{u}$, first sample policy $\pi\sim u$, and follow $\pi$ for the entire episode.  [^4] Note that we use a notation $\pi_{u}$ different from every-step mixing policy notation $\bar{\pi}_{u}$.
Importantly, even when the policies in $\Pi$ are all Markovian, $\pi_{u}$ may no longer be a Markovian policy: for example, the random draw of action $a_{2}$ does not only depend on the Markovian state $s_{2}$ but also the random policy $\pi$ drawn, which in turn may correlate with $a_{1}$.
To simplify the notations in our development below, following [80], we define causally conditional probabilities, when the agent uses Markovian policies or their first-step mixing:
Definition 14.
Given an MDP $\mathcal{M}$, the causally conditional probability of state sequence $s_{1:H}$ given action sequence $a_{1:H-1}$, is defined as:
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=\rho(s_{1})\prod_{h=1}^{H-1}P(s_{h+1}\mid s_{h},a_{h}).$
Given a (first-step mixture) of Markovian policy, its causally conditional probability of state sequence $a_{1:H}$ given action sequence $s_{1:H}$ is defined as:
  • For Markovian policy $\pi=\pi_{1:H}$, $\pi(a_{1:H}\parallel s_{1:H}):=\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$.
Note that $\pi(\cdot\|s_{1:H})$ and $\pi_{u}(\cdot\|s_{1:H})$ are valid probability distributions (e.g., $\sum_{a_{1:H}}\pi(a_{1:H}\|s_{1:H})=1$), however, the use of $\parallel$’ highlights its distinction from standard conditioning. For example, when executing Markovian policy $\pi$, the conditional probability of the actions given the states $\mathbb{P}^{\pi}(a_{1:H}|s_{1:H})\propto\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})\prod_{h=1}^{H-1}P(s_{h+1}\mid s_{h},a_{h})$, which is clearly different from its causally conditional counterpart.
We have the following (perhaps folklore) lemma for causally conditional probability (e.g. [80]).
@@ C.1 Preliminaries: First-step mixing and causally conditioned probabilities
For any Markovian policy $\pi$,
and for any first-step mixing of Markovian policies $\pi_{u}$,
@@
Eq. (4) follows by noting that both sides are equal to
$\rho(s_{1})\prod_{h=1}^{H-1}P(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
Eq. (5) follows by noting that both sides are equal to
$\sum_{\nu}u(\nu)\rho(s_{1})\prod_{h=1}^{H-1}P(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).\qed$
Lemma 16.
If $\bar{\pi}_{u}$ is an each-step policy in $\bar{\Pi}_{\mathcal{B}}$, then there exists some first-step mixture of policies in $\tilde{\mathcal{B}}$ that is equivalent to $\bar{\pi}_{u}$, i.e., they induce the same distribution over all length-$H$ trajectories.
Proof.
Define $\mu$ as a distribution over class $\tilde{\mathcal{B}}$ with $\mu(\nu):=\prod_{h=1}^{H}u(\nu_{h})$, for every $\nu=\nu_{1:H}$ in $\tilde{\mathcal{B}}$. Consider the joint action distribution under $\bar{\pi}_{u}$, which samples $\pi_{h}\sim u$ independently for each step and executes $a_{h}\sim\pi_{h}(\cdot\mid s_{h})$. The resulting causally conditional distribution over actions given state sequence $s_{1:H}$ is
$\bar{\pi}_{u}(a_{1:H}\|s_{1:H})=\prod_{h=1}^{H}\left(\sum_{\pi_{h}\in\mathcal{B}}u(\pi_{h})\,\pi_{h}(a_{h}\mid s_{h})\right).$
On the other hand, under the first-step mixture policy $\pi_{\mu}$ over $\tilde{\mathcal{B}}$, a full tuple $\nu=(\nu_{1},\dots,\nu_{H})$ is sampled once from $\mu$, and actions are drawn as $a_{h}\sim\nu_{h}(\cdot\mid s_{h})$. The resulting action distribution is
$\pi_{\mu}(a_{1:H}\|s_{1:H})=\sum_{\nu\in\tilde{\mathcal{B}}}\mu(\nu)\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$
@@
$\sum_{(\nu_{1},\dots,\nu_{H})\in\mathcal{B}^{H}}\left(\prod_{h=1}^{H}u(\nu_{h})\nu_{h}(a_{h}\mid s_{h})\right)=\prod_{h=1}^{H}\sum_{\pi_{h}\in\mathcal{B}}u(\nu_{h})\,\nu_{h}(a_{h}\mid s_{h}),$
@@
Therefore, $\bar{\pi}_{u}(a_{1:H}\|s_{1:H})=\pi_{\mu}(a_{1:H}\|s_{1:H}),$ and both policies induce the same trajectory distribution by Lemma 15. ∎
C.2 Proof of Theorem 6
Lemma 17.
Let $\tilde{\mathcal{B}}_{\mathrm{bc}}:=\{\pi=\pi_{1:H}\in\tilde{\mathcal{B}}:\pi_{h}(s_{h})=a_{h},\ \forall h\in[H],\forall(s,a)_{1:H}\in D_{\mathrm{off}}\}$ be the set of policies in $\tilde{\mathcal{B}}$ that agree with the expert on the offline dataset of $N_{\mathrm{off}}$ iid expert trajectories. Assume the expert policy $\pi^{E}$ is deterministic and realizable. Then, with probability at least $1-\delta$: for all $\pi$ in $\tilde{\mathcal{B}}_{\mathrm{bc}}$,
$J(\pi^{\mathrm{E}})-J(\pi)\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$ (6)
Consequently, for all $\pi^{n}$’s computed in Warm-Stagger (Algorithm 2), it holds that:
$J(\pi^{\mathrm{E}})-J(\pi^{n})\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$ (7)
Proof.
Eq. (6) is a direct consequence of behavior cloning’s guarantee applied to Markovian policy class $\tilde{\mathcal{B}}$ ([17, Corollary 2.1]). [^5] Our presentation of Theorem 2 uses a less general interpretation of that corollary, by restricting the policy class to be stationary.
Eq. (7) follows from Lemma 16 that $\mathbb{P}^{\pi^{i}}$ is a convex combination of $\mathbb{P}^{\pi}$’s for $\pi$’s in $\tilde{\mathcal{B}}_{\mathrm{bc}}$, as well as the fact that the expected return function $J(\pi)$ is linear in the trajectory distribution $\mathbb{P}^{\pi}$. ∎
Theorem 18 (Theorem 6 restated).
If Algorithm 2 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\mathrm{bc}}/\delta)}{N_{on}}\right)\right),$ (8)
Proof.